OpenStudy (anonymous):

To launch a communication satellite in a circular orbit 1000 km above the surface of the Earth, what orbital speed must be imparted to the satellite? Find the acceleration due to gravity at the orbit.

OpenStudy (anonymous):

equate a(due to g) to a=v^2/r a(due to g)=F(g)/m=|-GMm/r^2 x (1/m)| =GM/r^2 so v^2/r=GM/r^2, by this you can find velocity b/c r is known, and both the mass of Earth and the gravitational constants are known. Then you find the acceleration by plugging your velocity back to the equation a=v^2/r

OpenStudy (anonymous):

is r=1000km? @lailaiwd

OpenStudy (anonymous):

nope, you need to add the radius of Earth to 1000km

OpenStudy (anonymous):

to get r?

OpenStudy (anonymous):

I believe so. This is a common problem, and you can probably find similar solutions on Google.

OpenStudy (anonymous):

thanks a lot!! x

OpenStudy (anonymous):

definitely :D is that your picture? If so, you're pretty

OpenStudy (anonymous):

haha..yes it is... i get that doubt a lot!! lol

OpenStudy (anonymous):

v=5.98 x 10^24 m/s? a=2.20 m/s^2? @lailaiwd

OpenStudy (anonymous):

sorry... i'm not in the condition to to the numerics myself... it's almost 3 am, and I'm dozing

OpenStudy (anonymous):

its ok...maybe you should try get some sleep...